The new river gorge bridge in west virginia is a 516 m long steel arcch. How much will its length change between temperature extremes -18.6 degrees Celcius and 35.6 degrees celcius. Please show work so I can learn how to do this.
Answer #1

Coeeficient of thermal expansion for steel

alpha = 12*10^-6 /K

dT = 35.6- (-18.6) = 54.2

L0 = 516 m

L = L0(1 + alpha*dT)

L - L0 = L0*alpha*dT

dL = 516*12*10^-6*54.2

dL = 0.335 m